OpenStudy (anonymous):

sin(A+B) and sin(A-B) A=180 B=60 please help!

OpenStudy (anonymous):

I know the formulas, but don't know how to show the work or anything. pretty confusing.

OpenStudy (anonymous):

hey, A =180, B = 60 you don't know how many A - B is?

OpenStudy (anonymous):

I have to find those like first find sin(A+B) AND THEN sin (A-B)

OpenStudy (anonymous):

no need to put formula here, just A+ B and then A- B , take sin of the sum, or of the different

OpenStudy (anonymous):

the formula to solve those are sin(A+b)= sinAcosB+cosAsinB

OpenStudy (anonymous):

I know, but in case A, B are not special like that.

OpenStudy (anonymous):

but if you still want to take this, you know the formula, just plug them in the formula

OpenStudy (anonymous):

I have to find the exact value the only thing that throws me off is the 180

OpenStudy (anonymous):

ok, first, let my method gives you the right answer A + B = 240 , sin 240 =?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, your method, sin (A + B) = sinA cos B + sinB cos A yours is sin (180 + 60) = sin (180) cos (60) + sin (60) cos (180) sin 180 = 0 so, the first term =0 you just have sin (60) cos (180) sin (60)= sqr (3)/2. cos (180) =-1 so, sin (60) * c0s(180) = sqr (3)/2 ( -1 ) = - sqr (3) /2 \[\huge\color\red\IS~THAT~HARD?\]

OpenStudy (anonymous):

your'e the best!!!!!

OpenStudy (anonymous):

do the same with part B)

OpenStudy (anonymous):

I am gonna delete every thing, you are better do the stuff by yourself

OpenStudy (anonymous):

thanks for the help though.

OpenStudy (anonymous):

ok. but to me, just add A +B or A -B . A-B = 120, sin (120) = ???? quite easier than apply formula

OpenStudy (anonymous):

I have one question since my teacher never really teached us this how would you find the exact value of sin120?

OpenStudy (anonymous):

wait, I show you the link

OpenStudy (anonymous):

look up the values as you wish

OpenStudy (anonymous):

ooohh the circle !

OpenStudy (anonymous):

hehehe... did your prof show you that circle? I think he did

OpenStudy (anonymous):

Yes he did but very vaguely! but thank's to you you have brought back my memory :)

OpenStudy (anonymous):

OK ,